OpenStudy (anonymous):

Hi,I need your help with bisecting angles im stuck. Suppose JK bisects angle LJM. The measure of angle LJK = (-10x + 3) and the measure of KJM = (-x + 21)º. Which equation represents the best way to find the measure of angle LJM? (Hint: Sketch the situation first.)

OpenStudy (anonymous):

UM LJK=KJM?

OpenStudy (anonymous):



OpenStudy (anonymous):

hey... deja vu... you guys in the same class ?

OpenStudy (campbell_st):

as the angle LJM is bisected by JK the -10x + 3 = -x + 21 or -9x = 18 x = -2 substitute angle LJK = (-10 x -2) + 3 = 23 Angle KJM = -(-2) + 21 = 23 proving KJ bisects then LJM = 23 + 23 = 46